# GalaxyGenius: A Mock Galaxy Image Generator for Various Telescopes from Hydrodynamical Simulations

### v1

@@ 5 Summary
To adequately prepare for current and upcoming galaxy surveys in data processing and scientific analysis, we propose a Python package named GalaxyGenius, designed to generate galaxy images for various photometric surveys. It consists of three main modules (see Section 2.1): data preprocessing, ideal data cube generation, and mock observation. Specifically, as is shown in Section 2.2, the preprocessing module extracts properties of various particles from snapshots of hydrodynamical simulations and creates execution files for the following procedures. Subsequently, the data cube generation module assigns the particles with SEDs according to their types and properties and then performs the Monte Carlo radiative transfer process with SKIRT to generate IFU-like ideal data cubes as illustrated in Section 2.3. Finally, the mock observation module (see Section 2.4) constructs realistic galaxy images considering the throughputs, PSF and noise levels of a particular survey.
We present sample outputs of GalaxyGenius in Section 3, showcasing a suite of ongoing and planned photometric surveys conducted on both ground-based and space-based platforms, including CSST (Figure 3), Euclid (Figure 4), HST (Figure 5), JWST (Figure 6), Roman (Figure 7), and HSC (Figure 8), covering wavelengths from ultraviolet to infrared. The mock images are primarily based on the IllustrisTNG simulation. We also illustrate some mock images generated based on the EAGLE simulation (Figure 10), demonstrating the capability of GalaxyGenius to generate galaxy images for arbitrary simulations provided with necessary information for the preprocessing interfaces.
@@ Acknowledgements.
XCZ and NL acknowledge the support from The science research grants from the China Manned Space Project (No. CMS-CSST-2021-A01), and The CAS Project for Young Scientists in Basic Research (No. YSBR-062), and The Ministry of Science and Technology of China (No. 2020SKA0110100). H.Y. acknowledge the supports from the National Natural Science Foundation of China (Grant No. 11988101). Z.H. acknowledges support from the China Postdoctoral Science Foundation under Grant Number GZC20232990 and the National Natural Science Foundation of China (Grant No. 12403104). A.I acknowledges Postdoctoral Research Fund (R517L(N)) from Guangdong Technion Israel Institute of Technology and the Alliance of International Science Organizations (Grant No. ANSO-VF-2024-01). CW acknowledge the support from the China Manned Space Project (No. CMSCSST-2021-A03).References
@@ Appendix A Instrumental Parameters
Instrumental parameters employed to perform mock observations for selected filters illustrated in Section 3 for CSST, Euclid, HST WFC3, JWST NIRCam, Roman and HSC are outlined in Table 3, 4, 8, 6, 7 and 8 respectively. and demonstrate the wavelength range of individual filter. The sky backgrounds, are calculated using Equation 10 employing sky emission curves for each survey. and are dark current, readout, diameter of telescope and pixel scale of filters respectively. For CSST, we use tabular sky emission in https://hst-docs.stsci.edu/acsihb/chapter-9-exposure-time-calculations/9-7-tabular-sky-backgrounds. The instrumental parameters are obtained from private communications. Note that the number of exposure, , are determined by the focal plane design. While for Euclid, the emission curve is obtained from background model in https://irsa.ipac.caltech.edu/applications/BackgroundModel/ by specifying coordinates at and observing location at Earth-Sun L2 point. The instrumental parameters are obtained from https://sci.esa.int/web/euclid/-/euclid-vis-instrument and https://sci.esa.int/web/euclid/-/euclid-nisp-instrument for VIS and NISP respectively. For UVIS filters of HST WFC3, we employ the same emission curve as for CSST, while for IR filters, the emission curve is obtained from background model used for Euclid but with observing location changed to Earth orbit. The instrumental parameters for UVIS and IR are provided in https://www.stsci.edu/hst/instrumentation/wfc3/instrument-design/. The curve for JWST is acquired from JWST backgrounds tool (JBT) in https://github.com/spacetelescope/jwst_backgrounds using default parameters. And instrumental parameters are retrieved from https://jwst-docs.stsci.edu/jwst-near-infrared-camera/nircam-instrumentation/nircam-detector-overview/nircam-detector-performance#gsc.tab=0. For Roman, same background model used by Euclid is employed. We use the most pessimistic values for these instrumental parameters provided in https://roman.ipac.caltech.edu/sims/Param_db.html. While for ground-based HSC, we employ sky curve at Mauna Kea site given in https://www.gemini.edu/observing/telescopes-and-sites/sites#OptSky. The intrumental parameters employed are given in Aihara et al. (2018).
CSST | |||||||
Filters | |||||||
(Å) | 2560 | 3220 | 4070 | 5530 | 6940 | 8400 | 9250 |
(Å) | 3190 | 3990 | 5470 | 6860 | 8360 | 9950 | 10200 |
(e / s) | 0.0042 | 0.317 | 0.2803 | 0.3591 | 0.3708 | 0.2168 | 0.0638 |
(e / s) | 0.02 | ||||||
(e) | 5 | ||||||
(m) | 2 | ||||||
(arcsec) | 0.074 | ||||||
(s) | 150 | ||||||
4 | 2 | 2 | 2 | 2 | 2 | 4 |
Euclid | ||||
Categories | VIS | NISP | ||
Filters | VIS | Y | J | H |
(Å) | 5588 | 9630 | 11870 | 15460 |
(Å) | 8916 | 12000 | 15470 | 19960 |
(e / s) | 0.7162 | 3.6863 | 4.4556 | 3.8228 |
(e / s) | 0.005 | 0.02 | ||
(e) | 2.2 | 6 | ||
(m) | 1.2 | |||
(arcsec) | 0.1 | 0.3 | ||
(s) | 600 | |||
1 |
HST | |||||||
Categories | UVIS | IR | |||||
Filters | F275W | F390W | F814W | F105W | F110W | F125W | F160W |
(Å) | 2477 | 3451 | 7138 | 9222 | 9302 | 11143 | 14130 |
(Å) | 3003 | 4361 | 9324 | 11860 | 13770 | 13850 | 16710 |
(e / s) | 0.0004 | 0.0181 | 0.0562 | 1.9397 | 3.2075 | 1.8718 | 13786 |
(e / s) | 0.0008 | 0.048 | |||||
(e) | 3.1 | 21 | |||||
(m) | 2.4 | ||||||
(arcsec) | 0.04 | 0.13 | |||||
(s) | 600 | ||||||
1 |
JWST | ||||||
Categories | Short | Long | ||||
Filters | F070W | F150W | F200W | F182M | F356W | F444W |
(Å) | 6241 | 13477 | 17784 | 17335 | 31767 | 39320 |
(Å) | 7751 | 16547 | 22054 | 19575 | 39456 | 49374 |
(e / s) | 0.3307 | 0.4721 | 0.4326 | 0.2502 | 1.1707 | 3.2627 |
(e / s) | 0.0019 | 0.0342 | ||||
(e) | 15.77 | 13.25 | ||||
(m) | 6.5 | |||||
(arcsec) | 0.031 | 0.063 | ||||
(s) | 600 | |||||
1 |
Roman | |||||||
Filters | F062 | F087 | F106 | F129 | F158 | F184 | F213 |
(Å) | 5000 | 7700 | 9400 | 11450 | 14000 | 16950 | 19650 |
(Å) | 7500 | 9700 | 11800 | 14400 | 17550 | 19900 | 22900 |
(e / s) | 7.1684 | 6.6275 | 7.0874 | 7.1373 | 7.0349 | 4.1791 | 3.5150 |
(e / s) | 0.1 | ||||||
(e) | 20 | ||||||
(m) | 2.36 | ||||||
(arcsec) | 0.11 | ||||||
(s) | 600 | ||||||
1 |
HSC | |||||
Filters | |||||
(Å) | 4110 | 5550 | 7010 | 8540 | 9340 |
(Å) | 5400 | 6900 | 8420 | 9280 | 10360 |
(e / s) | 49.9830 | 79.8137 | 91.8011 | 107.2050 | 198.5708 |
(e / s) | 0 | ||||
(e) | 4.5 | ||||
(m) | 8.2 | ||||
(arcsec) | 0.168 | ||||
(s) | 600 | ||||
1 |
@@ Appendix C Noise level
Assuming that a given limiting magnitude corresponds to signal-to-noise ratio , we have:
@@
where , and are flux in A-1, zero-point and limiting magnitude respectively. Note that is already integrated over solid angle. is in , converted from by Equation 6 with pivot wavelength for bandpass:
Here we define gain factor as:
This factor is designed for conversion between electron counts and flux, and is only relevant to the observational instrument. Consequently, the electron counts can be written as:
Insert this equation to Equation 13, and then we have:
The noise level is the denominator of Equation 13. Therefore, for a source with intensity in unit of , the noise level in electron counts would be:
Note that the source intensity in is converted to electron counts by gain factor :
Alternatively, the noise level can be expressed in by:
Finally, the noise level can be incorporated into each pixel of the galaxy image by sampling a Gaussian distribution with a mean of 0 and a standard deviation of or , depending on the unit. This approach does not necessitate instrumental parameters such as dark current and readout noise , thereby offering a straightforward method to estimate noise levels when certain instrumental parameters are unavailable, particularly for planning instruments. In summary, this approach or the one mentioned in Section 2.4 can both be employed when adding noises to galaxy images, considering the availability of certain parameters.

### v2

@@ 5 Summary
To adequately prepare for current and upcoming galaxy surveys in data processing and scientific analysis, we propose a Python package, GalaxyGenius, designed to generate galaxy images for various photometric surveys. It consists of three main modules (see Section 2.1): data preprocessing, ideal data cube generation, and mock observation. Specifically, as shown in Section 2.2, the preprocessing module extracts properties of various particles from snapshots of hydrodynamical simulations and creates execution files for the following procedures. Subsequently, the data cube generation module assigns the particles with SEDs, according to their types and properties and then performs the MC radiative transfer process with SKIRT to generate IFU-like ideal data cubes (as illustrated in Section 2.3). Finally, the mock observation module (see Section 2.4) constructs realistic galaxy images considering the throughputs, PSF and noise levels of a particular survey.
We present sample outputs of GalaxyGenius in Section 3, showcasing a suite of ongoing and planned photometric surveys conducted on both ground-based and space-based platforms, including CSST (Figure 3), Euclid (Figure 4), HST (Figure 5), JWST (Figure 6), Roman (Figure 7), and HSC (Figure 8), covering wavelengths from UV to IR. The mock images are primarily based on the IllustrisTNG simulation. We also illustrate some mock images generated based on the EAGLE simulation (Figure 10), demonstrating the capability of GalaxyGenius to generate galaxy images for arbitrary simulations provided with necessary information for the preprocessing interfaces.
@@ Acknowledgements.
We appreciate helpful suggestions and comments from Maarten Baes, Qi Zeng, Andrea Gebek, Nick Andreads and members of SKIRT team. XCZ and NL acknowledge the support from the science research grants from the China Manned Space Project (No. CMS-CSST-2021-A01), the CAS Project for Young Scientists in Basic Research (No. YSBR-062), and the Ministry of Science and Technology of China (No. 2020SKA0110100). H.Y. acknowledges the support from the National Natural Science Foundation of China (Grant No. 11988101). Z.H. acknowledges support from the China Postdoctoral Science Foundation under Grant Number GZC20232990 and the National Natural Science Foundation of China (Grant No. 12403104). A.I. acknowledges the Postdoctoral Research Fund (R517L(N)) from Guangdong Technion Israel Institute of Technology and the Alliance of International Science Organizations (Grant No. ANSO-VF-2024-01). CW acknowledges the support from the China Manned Space Project (No. CMSCSST-2021-A03). SS thanks for research grants from the National Key R&D Program of China (No. 2022YFF0503402), the China Manned Space Project with No. CMS-CSST-2025-A07, the National Natural Science Foundation of China (No. 12141302), and Shanghai Academic/Technology Research Leader (22XD1404200).CMS-CSST-2025-A07, National Natural Science Foundation of China (No. 12141302) and Shanghai Academic/Technology Research Leader (22XD1404200).References
@@
- Gong et al. (2025) Gong, J.-Y., Lin, W., Tang, L., & Lan, Y. 2025, arXiv e-prints, arXiv:2504.18042
@@
- Massa et al. (2014) Massa, D., Fitzpatrick, E. L., & Gordon, K. D. 2014, in American Astronomical Society Meeting Abstracts, Vol. 223, American Astronomical Society Meeting Abstracts #223, 111.06
@@
- Meny et al. (2007) Meny, C., Gromov, V., Boudet, N., et al. 2007, A&A, 468, 171
- Misiriotis et al. (2006) Misiriotis, A., Xilouris, E. M., Papamastorakis, J., Boumis, P., & Goudis, C. D. 2006, A&A, 459, 113
@@
- Tang et al. (2018) Tang, L., Lin, W., Cui, W., et al. 2018, ApJ, 859, 85
- Tang et al. (2020) Tang, L., Lin, W., & Wang, Y. 2020, ApJ, 893, 87
- Tang et al. (2023) Tang, L., Lin, W., Wang, Y., Li, J., & Lan, Y. 2023, ApJ, 959, 104
- Tang et al. (2021) Tang, L., Lin, W., Wang, Y., & Napolitano, N. R. 2021, MNRAS, 508, 3321
@@
- Thomas et al. (2017) Thomas, R., Le Fèvre, O., Le Brun, V., et al. 2017, A&A, 597, A88
- Thomas et al. (2021) Thomas, R., Pentericci, L., Le Fèvre, O., et al. 2021, A&A, 650, A63
- Thomas et al. (2020) Thomas, R., Pentericci, L., Le Fevre, O., et al. 2020, A&A, 634, A110
@@
- Zhang et al. (2023) Zhang, R., Yuan, H., & Chen, B. 2023, ApJS, 269, 6
- Zhang & Green (2025) Zhang, X. & Green, G. M. 2025, Science, 387, 1209
@@ Appendix A Instrumental Parameters
Instrumental parameters employed to perform mock observations for selected filters illustrated in Section 3 for CSST, Euclid, HST WFC3, JWST NIRCam, Roman, and HSC are outlined in Figure 15. and demonstrate the wavelength range of individual filter. The sky backgrounds, are calculated using Equation 10 employing sky emission curves for each survey. and are dark current, readout, diameter of telescope and pixel scale of filters respectively. For CSST, we use tabular sky emission in https://hst-docs.stsci.edu/acsihb/chapter-9-exposure-time-calculations/9-7-tabular-sky-backgrounds. The instrumental parameters are obtained from private communications. Note that the number of exposure, , are determined by the focal plane design. While for Euclid, the emission curve is obtained from background model in https://irsa.ipac.caltech.edu/applications/BackgroundModel/ by specifying coordinates at and observing location at Earth-Sun L2 point. The instrumental parameters are obtained from https://sci.esa.int/web/euclid/-/euclid-vis-instrument and https://sci.esa.int/web/euclid/-/euclid-nisp-instrument for VIS and NISP respectively. For UVIS filters of HST WFC3, we employ the same emission curve as for CSST, while for IR filters, the emission curve is obtained from background model used for Euclid but with observing location changed to Earth orbit. The instrumental parameters for UVIS and IR are provided in https://www.stsci.edu/hst/instrumentation/wfc3/instrument-design/. The curve for JWST is acquired from JWST backgrounds tool (JBT) in https://github.com/spacetelescope/jwst_backgrounds using default parameters. And instrumental parameters are retrieved from https://jwst-docs.stsci.edu/jwst-near-infrared-camera/nircam-instrumentation/nircam-detector-overview/nircam-detector-performance#gsc.tab=0. For Roman, same background model used by Euclid is employed. We use the most pessimistic values for these instrumental parameters provided in https://roman.ipac.caltech.edu/sims/Param_db.html. While for ground-based HSC, we employ sky curve at Mauna Kea site given in https://www.gemini.edu/observing/telescopes-and-sites/sites#OptSky. The intrumental parameters employed are given in Aihara et al. (2018).
@@ Appendix C Noise level
Assuming that a given limiting magnitude corresponds to the S/N , we have
@@
where , and are flux in A-1, zero-point and limiting magnitude respectively. We note that is already integrated over solid angle. is in , converted from by Equation 6 with a pivot wavelength, , for a bandpass of
Here, we define gain factor as
This factor is designed for conversion between electron counts and flux, and is only relevant to the observational instrument. Consequently, the electron counts can be written as
Insert this equation to Equation 13 and then we have
The noise level is the denominator of Equation 13. Therefore, for a source with intensity, , in unit of , the noise level, , in electron counts would be:
We note that the source intensity in is converted to electron counts by gain factor :
Alternatively, the noise level can be expressed in via
Finally, the noise level can be incorporated into each pixel of the galaxy image by sampling a Gaussian distribution with a mean of 0 and a standard deviation of or , depending on the unit. This approach does not necessitate instrumental parameters such as dark current, , and readout noise, , thereby offering a straightforward method to estimate noise levels when certain instrumental parameters are unavailable, particularly for planning instruments. In summary, this approach (or the one mentioned in Section 2.4) can both be employed when adding noise to galaxy images, considering the availability of certain parameters.